OpenStudy (anonymous):

a) Find the product of the 2 matrices below showing all your working. [2 7] [3 8] [-1 4] [5 0] b) Find the product of the 2 matrices below showing all your working and explain the significance of this result. [3 8] [2 7] [5 0] [-1 4] If A= [7 8] [5 2] Find A^-1

OpenStudy (anonymous):

hm.. how to do this one..?

OpenStudy (amistre64):

i believe its... 2.3+7.5 2.8+7.0 -1.3+4.5 -1.8+4.0

OpenStudy (anonymous):

hm.. I don't know whether it's right or wrong..

OpenStudy (amistre64):

its right for a

OpenStudy (amistre64):

2.3 means 2(3)... just in a lazy format

OpenStudy (anonymous):

its like this: (2x3)+(7x5) (8x2)+(7x0) (-1x3)+(4x5) (-1x8)+(4x0)

OpenStudy (amistre64):

6+35 = 41 16+0 = 16 -3+20=17 -8+0=-8

OpenStudy (anonymous):

Yup, I got the answer for a.. but, I'm stuck in b and c...

OpenStudy (anonymous):

do the same process in b

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how to explain the significance of this result??

OpenStudy (amistre64):

b is the same process: [3 8] [2 7] [5 0] [-1 4] 3.2+8.-1 = -2 3.7+8.4= 53 5.2+0.-1= 10 5.7+0.4=35

OpenStudy (amistre64):

c; find the inverse... [7 8] 7.a+8.c=1 7.b+8.d=0 [5 2] 5.a+2.c=0 5.b+2.d=1

OpenStudy (anonymous):

in part c its doing the nverse, do you know the proces of doing it?

OpenStudy (amistre64):

[a b] [c d] will be the inverse

OpenStudy (amistre64):

the short cut is to flip and negate the 7 and 2 :)

OpenStudy (amistre64):

it is the 7 and 2 right?

OpenStudy (amistre64):

flip a d and negate b c

OpenStudy (amistre64):

[2 -8] [-5 7]

OpenStudy (anonymous):

ok I've flipped a d and negate b c..

OpenStudy (amistre64):

that is the inverse of A :)

OpenStudy (anonymous):

ok thank you..! So after inverse, there's nothing else right?

OpenStudy (anonymous):

yup amistre got it,, theres nothing else

OpenStudy (anonymous):

Thanks so much guys!

OpenStudy (anonymous):

wc